INAF queue observing, 2016 November 5 - 6

Observer: F. Cusano, A. Rossi
Telescope Operator: D. Gonzalez
Support Astronomer: O. Kuhn on call
Instrument: LBC

Summary

At the sunset the sky is covered by clouds, humidity at 100%. At 3:40 the sky is clear and humidity is around 80%.
The TO opens the dome at 3:55. LUCI2 still not available
We use LBCs.

Night duration (12\xB0 el twilight) 18:30-5:30 ( 01:30-12.30 UT) = 11h 00m

Observing time: h m (?%)

Weather time loss: h (?%)

technical time loss h (?%)

Data Summary

Project_RUN INST Objects Filter ExpTime Notes
Standard LBC Sa92 S,g,r -V,i,z   To complete Baglio ToO
Annibali R45 LBC UGC12613 g,r 40min UGC12613_3.ob to complet the program
Annibali R2 LBC UGC1249 g,r 1h completed
Castellano A LBC MACSJ0717 Std_U,V 1.7h  
Annibali R8 LBC UGC3755 g,r 20min completed

Details

3:55 The TO opens the dome

4:07 sending copointing of Baglio
  • 4:11 dohybrid, did not converge; 4:21 run again dohybrid
  • stop again dohybrid can not find pupils in the blue channel, 4:28 we run dofpia, X2
  • 4:33 blue not converging we try to change field we move to standard field SA92
  • 4:35 the TO and Olga decide to reset LBC
  • 4:48 dohybrid

5:12 observing standard SA92 (filter: Sgr - Viz) * seeing 0.8"

5:22 slew to Baglio field (2 minute exp)
  • stars elongated, 5:31 we run dofpia again
  • 5:35 Baglio field again, stars look good now, seeing 0.6"

5:40 slewing to copointing of Annibali R45 (UGC12613)
  • dofpia, X2

5:53 starting UGC12613_1.ob, Annibali R45
  • 5:55 seeing 0.7"
  • 6:10 seeing 1" on images

6:19 UGC12613_calib.ob
  • seeing 0.9" on images, dimm seeing 0.8"

6:23 dofpia

6 28 UGC12613_2.ob
  • dimm seeing 0.8"
  • seeing 0.7" on images at 6:35

6:53 UGC12613_3.ob
  • stars elongated in the blue channel, seeing 1.1" out of spec, almost reached the limit in airmass, we change target

7:01 slewing to Annibali R2 (UGC1249)
  • UGC1249_point.ob
  • dofpia
  • lbcrangebal

7:13 executing UGC1249_1.ob
  • dimm seeing 0.9", on images 0.7"
  • seeing 0.9" at 7:19
  • seeing 0.7" at 7:35

7:40 dofpia

7:45 executing UGC1249_2.ob
  • seeing 0.9" at 7:55
  • seeing 0.7" 8:06

8:13 executing UGC1249_3.ob
  • seeing 0.9" at 8:40

8:44 slewing to Castellano Run A, copointing
  • dofpia
  • lbcrangebal (crashed than restored)
  • seeing 0.8"

8:56 executing Castellano Run A, Castellano_MACSJ0717_UV_14.ob
  • 9:02 seeing 0.85" on images
  • 9:12 seeing 0.9"
  • 9:20 seeing 0.9"

9:37 dofpia

9:41 Castellano_MACSJ0717_UV_15.ob
  • seeing 0.9" at 9:50
  • seeing 0.7" at 10:03

10:23 Castellano_MACSJ0717_UV_16.ob
  • seeing 0.7" at 10:33
  • seeing 0.9" at 10:45
  • seeing 0.7" at 10:59

11:06 slewing to Annibali R8
  • dofpia

11:16 executing UGC3755_3.ob (Annibali R8)
  • seeing 0.7" at 11:25
  • seeing 0.6" at 11:37

11:41 slewing to Castellano Run A,
  • dofpia
  • 11:45 the humidity is climbing, now at 87%

11:50 Castellano_MACSJ0717_UV_17.ob
  • seeing 0.8" at 11:59
  • 12:03 humidity stable at 89%
  • seeing 0.7" at 12:07
  • 12:08 the telescope stop tracking, we have to send the ob again

12:11 executing again Castellano_MACSJ0717_UV_17.ob after telescope problem
  • stop the ob, the humidity is 92% and rising

-- %USERSIG{FeliceCusano - 2016-11-06}%

Comments

 
Topic revision: r5 - 07 Dec 2016, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback